how do i find the value of "y"
-7y - 2 = 68

Answer by snowwolf11(1)

You must isolate the variable "y" to be by itself on one side of the '=' sign, and then the answer is on the other side of the '=' sign.
To isolate "y", you must reverse the order of operations. First, ask yourself "what is being added or subtracted to the variable 'y'?". The answer is: "2 is being subtracted". You undo this subtraction by adding "2" to each side of the equation (whatever you add to one side of the equation, you must add the same to the other side).
This yields the following modified equation:
-7y = 68 + 2
Combining terms on the right side of the equation gives:
-7y = 70
Now, to finish isolating "y" we must undo the multiplication by "-7". This multiplication can be undone by division. So, we divide both sides by "-7", giving:
-7y/-7 = 70/-7
The multiplication and division by -7 on the left side of the equation will cancel out, leaving the following:
y = - 70/7
Since 70/7 = 10, the answer can be simplified:
y = -10